OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f + g)(−4)?

OpenStudy (mathstudent55):

(f + g)(x) = f(x) + g(x) Evaluate each function at -4 and add the results.

OpenStudy (anonymous):

I'm sorry but I don't get what that means. Sorry I am quite bad at Algebra, Geometry is my thing.

OpenStudy (mathstudent55):

To evaluate f(-4), do this: f(-4) = 3(-4) - 2 Can you finish that now? Then do the same for g(-4): g(-4) = (-4)^2 - 1 Can you finish that also? Then just add what you get for f(-4) and g(-4)

OpenStudy (anonymous):

Oh alrighty, so its 1?

OpenStudy (mathstudent55):

f(-4) = 3(-4) - 2 = -12 - 2 = -14 g(-4) = (-4)^2 - 1 = 16 - 1 = 15 f(-4) + g(-4) = -14 + 15 = 1 You are correct.

OpenStudy (anonymous):

Awesome thank you! One final question, when for example the equation has f[g(5)] where it is multiplied how do you solve that?

OpenStudy (mathstudent55):

This is not multiplied. This is first finding the value of g(5), then inserting that value into f(x).

OpenStudy (anonymous):

Oh alright that just saved me from getting multiple questions wrong.

OpenStudy (mathstudent55):

The way to show a product of functions is \( (f \cdot g)(x) = f(x) \cdot g(x)\) The following is the composition of functions, which is what you had above: \( (f \circ g)(x) = f(g(x)) \) The line above is read: The composition of function f and function g is function f evaluated at g.